# A Classification-Driven Likelihood Ratio Method for Familial DNA Testing

### v1

@@ 2.1 Existing Test Statistics
Consider a population divided into $R>1$ distinct subpopulations, each with its own allele frequency distribution, denoted as $f_{i}$ for the $i^{\text{th}}$ subpopulation. Below, we define four existing methods for conducting familial DNA testing when the profiles $X_{1}$ and $X_{2}$ are given.
@@ 2.2 Proposed Test Statistic
To understand how Naive Bayes classification operates, consider a population divided into $R>1$ distinct subpopulations. Given a DNA profile $B$, the task is to determine which subpopulation it belongs to. This can be achieved by calculating the probability that the profile $B$ originates from each allele frequency distribution. Mathematically, let $A_{i}$ be the event such that a DNA profile $B$ corresponds to the $i^{\text{th}}$ subpopulation. By using the Bayes’ rule, for i{1, 2,,R},i\ $\in\{$1,\ 2,\ $\ldots$,\ R\}, we have
@@
On the right-hand side of the equation, the subpopulation prior probability $P\left(A_{i}\right)$ can be determined using the proportion of each subpopulation. The term $P\left(B\mid A_{i}\right),$ which represents the probability of observing a DNA profile $B$ given that the associated subpopulation is $i,$ can be computed using allele frequency data from the $i^{\text{th}}$ subpopulation. An illustrative example is presented here: consider a scenario with $R=4$ subpopulations and three independent loci, namely TPOX, vWA, and D13S317. The DNA profiles of two individuals, $X_{1}$ and $X_{2}$, are presented in Table 1:
@@
Locus
DNA profiles TPOX vWA D13S317
@@
The calculations for other values of $k$ and $i$ can be performed in the same manner. When all values are combined, the results in this context are $P\left(A_{1}\mid X_{1},X_{2}\right)=0.2903,P\left(A_{2}\mid X_{1},X_{2}\right)=0.0431,P\left(A_{3}\mid X_{1},X_{2}\right)=0.0729,$ and $P\left(A_{4}\mid X_{1},X_{2}\right)=0.5938$.
@@
The following provides the formal definition of the proposed statistic. Given DNA profiles for two individuals, we aim to simultaneously classify which subpopulation these profiles belong to. In the first step, Naive Bayes classification is applied. Then, the likelihood ratio statistic is used within a single population framework, based on the classification result. This statistic, denoted as LRCLASS, is mathematically expressed as:
@@ 2.3 Algorithms for Power Comparison
Recall that power refers to the probability of correctly rejecting the null hypothesis when the alternative hypothesis is true. To determine this, we first generate a null distribution to establish a criterion for rejecting the hypothesis, and then create the alternative hypothesis to estimate the power. The algorithm is outlined below.
@@
    Simulate a pair of unrelated DNA profiles using $m=15$ loci. The subpopulation information of the individuals is randomly selected from a multinomial distribution with success probabilities $p_{NO}=0.11083,\,p_{NE}=0.36944,\,p_{CT}=0.35383,\,p_{SO}=0.16590$. These numbers are estimated from Thai population database (2023).
@@
Once the power for each statistic is obtained, we aim to compare them at the same false positive rate (FPR) $\alpha$. To determine an appropriate value for $\alpha$, we can refer to previous studies in the Colorado prison population, which suggested that for parent-child and full-sibling tests, $\alpha$ should be $1.7\times 10^{-5}$ and $1.2\times 10^{-5},$ respectively (Kooakachai et al. 2019).
@@ 2.4 Classification Comparison
We have developed the LRCLASS statistic, which classifies two DNA profiles into a single subpopulation simultaneously. The classification method employed was the Naive Bayes classifier. However, in statistics and machine learning, numerous other classification techniques are available. In this section, we describe an alternative classification method—logistic regression—and discuss how its performance compares to the previously proposed Naive Bayes classifier.
@@ 3.1 Power Comparison
[…]
Subsequently, Table 4 shows the power of each test for the parent-child relationship at an FPR of $1.7\times 10^{-5}$ and for the full-sibling relationship at an FPR of $1.2\times 10^{-5}$. Additionally, the table includes 95% confidence intervals for the power estimates to provide a more robust measure of precision.
@@
To further investigate the performance of LRCLASS, power comparisons were visualized by plotting its power against each of the existing methods. The analysis was extended to include a range of FPRs, spanning from $10^{-6}$ to $4\times 10^{-5}$. From Figure 1, the results indicate that LRCLASS consistently outperforms LRMAX across nearly the entire range of FPRs considered. However, its performance relative to LRLAF, LRAVG, and LRMIN is less conclusive, as no clear dominance is observed throughout the FPR range analyzed. For the full-sibling relationship test, as shown in Figure 2, LRCLASS demonstrates superior performance compared to all other statistics across the entire range of FPRs considered.
@@ 3.2 Classification Comparison
Tables 5 and 6 present the performance evaluation of each classification method using the validation dataset and the test dataset, respectively. The best-performing method for each criterion is emphasized in bold. It is important to note that the criteria listed in the first column of Tables 5 and 6 represent technical terms, which are explained here:
@@
Logistic Mthd A Logistic Mthd B Naive Bayes
Precision 0.6975 0.6919 0.6916
Recall 0.6985 0.6944 0.6938
Specificity 0.8500 0.8488 0.8511
Accuracy 0.8259 0.8237 0.8232
Error 0.1741 0.1763 0.1768
F1-score 0.6864 0.6824 0.6857
[…]
@@
Logistic Mthd A Logistic Mthd B Naive Bayes
Precision 0.5003 0.4928 0.4924
Recall 0.4622 0.4584 0.4693
Specificity 0.7915 0.7896 0.7869
Accuracy 0.7219 0.7197 0.7237
Error 0.2781 0.2803 0.2763
F1-score 0.4435 0.4378 0.4462
[…]
First, we recall four standard terms: True Positive (TP) refers to a sample correctly identified as positive, while True Negative (TN) is a sample correctly identified as negative. False Positive (FP) occurs when a sample is incorrectly classified as positive, and False Negative (FN) occurs when a sample is incorrectly classified as negative. The metrics for classification are defined as follows:
  • Precision: the number of samples correctly classified as positive out of all samples predicted to be positive by the model. Mathematically, it is TP/(TP+FP). Since there are four subpopulations, each is treated as the positive class in turn, and precision is calculated for each. The average precision across the four subpopulations is then used as the final measure. This averaging approach is applied consistently to all other measures.
  • Recall: the number of samples correctly predicted to belong to the positive class out of all samples that actually belong to the positive class, written as TP/(TP+FN).
  • Specificity: the number of samples correctly predicted to be in the negative class out of all samples that actually belong to the negative class, given by TN/(FP+TN).
  • Accuracy: the number of samples correctly classified out of all samples in the test set, calculated as (TP+TN)/(TP+FP+TN+FN).
  • Error: the number of samples incorrectly classified out of all samples in the test set, given by (FP+FN)/(TP+FP+TN+FN).
[…]
@@ 4 Discussions
In terms of power comparison, as shown in Table 4, the performance can be categorized into two distinct groups. The first group includes LRLAF, LRMIN, and LRCLASS, all of which demonstrate strong performance in both the parent-child and full-sibling tests. The second group consists of LRAVG and LRMAX, which clearly exhibit lower power compared to the other statistics. This observation is consistent with the results of previous studies conducted on the Colorado population (Kooakachai et al. 2019). In particular, we recommend using the proposed statistic LRCLASS for the full-sibling test in the Thai population, as it consistently outperforms all other statistics, as illustrated in Figure 2. There are no significant limitations associated with using these test statistics, as the computational time required for each method is comparable.
@@ 5 Conclusions
6 List of abbreviations
DNA Deoxyribonucleic Acid
LR Likelihood ratio
LRLAF Likelihood ratio statistic involving the use of average allele frequency distribution
LRMAX Likelihood ratio statistic based on the maximum of all available likelihood ratios
LRMIN Likelihood ratio statistic based on the minimum of all available likelihood ratios
LRAVG Likelihood ratio statistic based on the average of all available likelihood ratios
LRCLASS Likelihood ratio statistic involving a classification step
Y-STR Short tandem repeat on the Y-chromosome
STR Short tandem repeat
USA United States of America
A Adenine
T Thymine
G Guanine
C Cytosine
HVR1 Hypervariable region 1
CIFS Central Institute of Forensic Science
NO North
NE Northeastern
CT Central
SO South
IBD Identical by descent
Mb Megabase
prob Probability
subpop Subpopulation
FPR False positive rate
PC Parent-child
SB Full-sibling
CI Confidence interval
TP True positive
TN True negative
FP False positive
FN False negative
Mthd Method
Pred Predict
7 Declarations
@@
The training datasets generated during the current study, along with the test dataset, are available from the corresponding author upon reasonable request. The allele frequency distributions, which serve as the basis for the generated datasets, can be found in the public literature (Shotivaranon et al. 2009).
Competing interests
The authors declare that they have no competing interests.
Funding
This research did not receive any financial support from institutions or organizations.

### v2

@@ 2.3 Existing Test Statistics
Consider a population divided into $R>1$ distinct subpopulations, each with its own allele frequency distribution, denoted as $f_{i}$ for the $i^{\text{th}}$ subpopulation. We define four existing methods for conducting familial DNA testing when the profiles $X_{1}$ and $X_{2}$ are given.
@@ 2.4 Proposed Test Statistic
Recall that we have a population divided into $R>1$ distinct subpopulations. Given a DNA profile $B$, the task is to determine which subpopulation it belongs to. Let $A_{i}$ be the event such that a DNA profile $B$ corresponds to the $i^{\text{th}}$ subpopulation. By using the Bayes’ rule, for i{1, 2,,R},i\ $\in\{$1,\ 2,\ $\ldots$,\ R\}, we have
@@
On the right-hand side of the equation, the subpopulation prior probability $P\left(A_{i}\right)$ can be determined using the proportion of each subpopulation. The term $P\left(B\mid A_{i}\right),$ which represents the probability of observing a DNA profile $B$ given that the associated subpopulation is $i,$ can be computed using allele frequency data from the $i^{\text{th}}$ subpopulation. An illustrative example is presented here: consider a scenario with $R=4$ subpopulations and three independent loci, namely TPOX, vWA, and D13S317. The DNA profiles of two individuals, $X_{1}$ and $X_{2}$, are presented in Table 1.
@@
    DNA profiles     Locus
    TPOX     vWA     D13S317
@@
The calculations for other values of $k$ and $i$ can be performed in the same manner. When all values are combined, the results in this context are as follows: $P\left(A_{1}\mid X_{1},X_{2}\right)=0.2903,P\left(A_{2}\mid X_{1},X_{2}\right)=0.0431,P\left(A_{3}\mid X_{1},X_{2}\right)=0.0729,$ and $P\left(A_{4}\mid X_{1},X_{2}\right)=0.5938$.
@@
Next, we formally define the proposed statistic. Given the DNA profiles of two individuals, we first use a Naive Bayes classifier to jointly determine their most likely subpopulation. Once the classification is obtained, the likelihood ratio is computed under a single-population framework corresponding to the assigned subpopulation. This statistic, denoted as LRCLASS, is mathematically expressed as:
@@ 3.1 Algorithms for Power Comparison
Recall that power refers to the probability of correctly rejecting the null hypothesis when the alternative hypothesis is true. To determine this, we first generate a null distribution to establish a criterion for rejecting the null hypothesis, and then create the alternative hypothesis to estimate the power. The algorithm is outlined below.
@@
    Simulate a pair of unrelated DNA profiles using $m=15$ loci. The subpopulation information of the individuals is randomly selected from a multinomial distribution with success probabilities $p_{NO}=0.11083,\,p_{NE}=0.36944,\,p_{CT}=0.35383,\,p_{SO}=0.16590$. These estimates are derived from the Thai population database provided by the National Statistical Office of Thailand (2023).
@@
Once the power for each statistic is obtained, we aim to compare them at the same false positive rate (FPR) $\alpha$. To determine an appropriate value for $\alpha$, we can refer to previous studies in the Colorado prison population, which suggested that for parent-child and full-sibling tests, $\alpha$ should be $1.7\times 10^{-5}$ and $1.2\times 10^{-5},$ respectively (Kooakachai et al., 2019).
@@ 3.2 Classification Comparison
We have developed the LRCLASS statistic, which classifies two DNA profiles into a single subpopulation simultaneously. The classification method employed was the Naive Bayes classifier. However, in statistics and machine learning, numerous other classification techniques are available. In this section, we describe an alternative classification method—multinomial logistic regression—and discuss how its performance compares to the previously proposed Naive Bayes classifier.
@@
3.3 Evaluation Metrics for Multi-class Classification
We begin by reviewing four standard classification terms. A True Positive (TP) is a sample correctly identified as positive, while a True Negative (TN) is correctly identified as negative. A False Positive (FP) occurs when a negative sample is incorrectly labeled as positive, and a False Negative (FN) occurs when a positive sample is incorrectly labeled as negative. The corresponding classification metrics are defined as follows:
  • Accuracy: the number of samples correctly classified out of all samples in the test set, calculated as
    $\text{Accuracy}=\frac{\text{TP}+\text{TN}}{\text{TP}+\text{TN}+\text{FP}+\text{FN}}\in[0,1].$
  • Precision: the number of samples correctly classified as positive out of all samples predicted to be positive by the model. Mathematically, it is
    $\text{Precision}=\frac{\text{TP}}{\text{TP}+\text{FP}}\in[0,1].$
  • Recall: the number of samples correctly predicted to belong to the positive class out of all samples that actually belong to the positive class, written as
    $\text{Recall}=\frac{\text{TP}}{\text{TP}+\text{FN}}\in[0,1].$
  • Specificity: the number of samples correctly predicted to be in the negative class out of all samples that actually belong to the negative class, given by
    $\text{Specificity}=\frac{\text{TN}}{\text{TN}+\text{FP}}\in[0,1].$
  • Error: the number of samples incorrectly classified out of all samples in the test set, given by
    $\text{Error}=\frac{\text{FP}+\text{FN}}{\text{TP}+\text{TN}+\text{FP}+\text{FN}}\in[0,1].$
  • F1-score: the harmonic mean of precision and recall for the positive class, defined as
    $\text{F1-score}=\frac{2\times\text{Precision}\times\text{Recall}}{\text{Precision}+\text{Recall}}\in[0,1].$
For a classification task with $n$ samples distributed among $R\geq 3$ classes, the classifier’s performance can be summarized using an $R\times R$ confusion matrix. Each entry $n_{ij}$ at the intersection of row $i$ and column $j$, for $i,j=1,\ldots,R$, represents the number of samples truly belonging to class $i$ but predicted as class $j$. The same evaluation metrics used in binary classification can be extended to this multiclass setting.
Two common strategies can be used to compute evaluation metrics (Rainio et al., 2024). The first constructs a separate $2\times 2$ confusion matrix for each of the $R$ classes, with components defined as:
$\mathrm{TP}_{i}=n_{ii},\quad\mathrm{TN}_{i}=\sum_{j\neq i}\sum_{k\neq i}n_{jk},\quad\mathrm{FN}_{i}=\sum_{j\neq i}n_{ij},\quad\mathrm{FP}_{i}=\sum_{j\neq i}n_{ji},$
[…]
The next two metrics are Cohen’s Kappa $(\kappa)$ and the Matthews Correlation Coefficient (MCC), as presented in Rainio et al. (2024). Both measures offer advantages over basic evaluation metrics in multi-class settings, particularly when dealing with imbalanced class distributions. Cohen’s Kappa incorporates an adjustment for agreement occurring by chance, making it more dependable when class proportions differ substantially. MCC, on the other hand, accounts for all entries in the confusion matrix and yields a more balanced, comprehensive assessment of classifier performance. Their formulas are given as follows:
$\kappa$ $=\frac{p_{0}-p_{e}}{1-p_{e}}\quad\text{with}\quad p_{0}=\frac{1}{n}\sum_{i=1}^{R}n_{ii},\quad p_{e}=\frac{1}{n^{2}}\sum_{i=1}^{R}n_{i\cdot}n_{\cdot i},$
MCC $=\frac{n\sum_{i=1}^{R}n_{ii}-\sum_{i=1}^{R}n_{i\cdot}n_{\cdot i}}{\sqrt{(n^{2}-\sum_{i=1}^{R}n_{i\cdot}^{2})(n^{2}-\sum_{i=1}^{R}n_{\cdot i}^{2})}},$
where $n_{i\cdot}=\sum_{j=1}^{R}n_{ij}$, $n_{\cdot i}=\sum_{j=1}^{R}n_{ji}$ and $n=\sum_{i=1}^{R}\sum_{j=1}^{R}n_{ij}$. We will evaluate and compare the classification performance using the metrics outlined above.
@@ 4.1 Power Comparison
Subsequently, Table 4 shows the power of each test for the parent-child relationship at an FPR of $1.7\times 10^{-5}$ and for the full-sibling relationship at an FPR of $1.2\times 10^{-5}$. Additionally, the table includes 95% confidence intervals for the power estimates, calculated using the exact Binomial test, to provide a robust measure of precision.
@@
To further investigate the performance of LRCLASS, power comparisons were visualized by plotting its power against each of the existing methods. The analysis was extended to include a range of FPRs, spanning from $10^{-6}$ to $4\times 10^{-5}$. Figures 1 and 2 depict the power curves for parent-child and full-sibling relationship testing, respectively, as functions of the FPR. In these figures, power is displayed on the y-axis, with the statistic corresponding to the upper curve demonstrating greater power than that associated with the lower curve. Additionally, the vertical lines in Figures 1 and 2 represent the suggested FPRs of $1.7\times 10^{-5}$ and $1.2\times 10^{-5},$ respectively.
From Figure 1, the results indicate that LRCLASS consistently outperforms LRMAX across nearly the entire range of FPRs considered. However, its performance relative to LRLAF, LRAVG, and LRMIN is less conclusive, as no clear dominance is observed throughout the FPR range analyzed. For the full-sibling relationship test, as shown in Figure 2, LRCLASS demonstrates superior performance compared to all other statistics across the entire range of FPRs considered.
@@ 4.2 Classification Comparison
Tables 5 and 6 present the performance evaluation of each classification method using the validation dataset and the test dataset, respectively. The metrics we used include micro- (Mic.) and macro- (Mac.) averaged accuracy, precision, recall, specificity, error rate, and F1-score, along with Cohen’s Kappa and Matthews Correlation Coefficient (MCC). The best-performing method for each criterion is emphasized in bold.
@@
  Metrics   Naive Bayes   Logistic Mthd A   Logistic Mthd B
  Accuracy   Mac.   0.8232   0.8259   0.8237
  Mic.   0.8469   0.8493   0.8472
  Precision   Mac.   0.6916   0.6975   0.6919
  Mic.   0.6938   0.6985   0.6944
  Recall   Mac.   0.6938   0.6985   0.6944
  Mic.   0.6938   0.6985   0.6944
  Specificity   Mac.   0.8511   0.8500   0.8488
  Mic.   0.8979   0.8995   0.8981
  Error   Mac.   0.1768   0.1741   0.1763
  Mic.   0.1531   0.1507   0.1528
  F1-score   Mac.   0.6857   0.6864   0.6824
  Mic.   0.6938   0.6985   0.6944
  Cohen’s Kappa   0.5504   0.5548   0.5490
  MCC   0.5537   0.5605   0.5542
@@
  Metrics   Naive Bayes   Logistic Mthd A   Logistic Mthd B
  Accuracy   Mac.   0.7237   0.7219   0.7197
  Mic.   0.7346   0.7311   0.7292
  Precision   Mac.   0.4924   0.5003   0.4928
  Mic.   0.4693   0.4622   0.4584
  Recall   Mac.   0.4693   0.4622   0.4584
  Mic.   0.4693   0.4622   0.4584
  Specificity   Mac.   0.7869   0.7915   0.7896
  Mic.   0.8231   0.8207   0.8195
  Error   Mac.   0.2763   0.2781   0.2803
  Mic.   0.2654   0.2689   0.2708
  F1-score   Mac.   0.4462   0.4435   0.4378
  Mic.   0.4693   0.4622   0.4584
  Cohen’s Kappa   0.2607   0.2566   0.2508
  MCC   0.2765   0.2693   0.2636
@@ 5 Discussions
In terms of power comparison, as shown in Table 4, the performance can be categorized into two distinct groups. The first group includes LRLAF, LRMIN, and LRCLASS, all of which demonstrate strong performance in both the parent-child and full-sibling tests. The second group consists of LRAVG and LRMAX, which clearly exhibit lower power compared to the other statistics. This observation is consistent with the results of previous studies conducted on the Colorado population (Kooakachai et al., 2019). In particular, we recommend using the proposed statistic LRCLASS for the full-sibling test in the Thai population, as it consistently outperforms all other statistics, as illustrated in Figure 2. There are no significant limitations associated with using these test statistics, as the computational time required for each method is comparable.
@@ 6 Conclusions
Acknowledgement
@@
The training datasets generated during the current study, along with the test dataset, are available from the corresponding author upon reasonable request. The allele frequency distributions, which serve as the basis for the generated datasets, can be found in the public literature (Shotivaranon et al., 2009).